# Products of hyperbolic spaces

### v1

@@ 1 Introduction
If $X$ is a metric space possessing enough convexity structure, and $T:X\to X$ is a nonexpansive mapping, one can define, for any $x\in X$, the point $x_{t}$, called a ‘resolvent’, by the implicit definition $x_{t}=tx_{t}+(1-t)x$ (existence and uniqueness being guaranteed by the Banach contraction principle). One may then ask about the strong convergence of the path $(x_{t})$ (for $t\to 1$), and, for $X$ being a bounded convex subset of a Hilbert space, this was first proven in 1967, independently by Browder [5] and Halpern [8]. A little more than ten years later, the corresponding result for uniformly smooth Banach spaces, requiring significantly new ideas, was proven by Reich [20].
@@
The analysis of the Browder-Halpern result was obtained by Kohlenbach in 2011 [11]; its sequel, the work on Reich’s theorem, had to wait, like the original theorem itself, a decade until it was finally achieved by Kohlenbach and the author [14], being one of the most intricate results ever achieved by the proof mining program (see also the author’s recent extension [23] to Reich’s original setting of accretive operators). The goals of these endeavours were to obtain quantitative results, namely so-called ‘rates of metastability’; what is more relevant here is that a proof-theoretic analysis usually strips down a proof to its bare essence, leaving it more suitable to generalization.
Such a generalization was needed, as per Pinto [18], since the theorem had also been proven, in the meantime, in the setting of CAT(0) spaces by Saejung [21] (and then proof-theoretically analyzed by Kohlenbach and Leuştean [13]), and, thus, one had to have a framework to encompass both CAT(0) spaces and uniformly smooth Banach spaces, this being achieved in the class of uniformly smooth hyperbolic spaces. In addition, what Kohlenbach and the author had done to tame the proof of Reich’s theorem and make it amenable to analysis and to generalization was to introduce the additional uniform convexity hypothesis on the Banach space (thus, still maintaining $L^{p}$ spaces as a particular case of the theorem), a notion which had already been formulated in the hyperbolic setting by Leuştean [16, 17] in the form of $UCW$-hyperbolic spaces. Thus, Pinto was able to show that Reich’s theorem and its consequences do hold for uniformly smooth $UCW$-hyperbolic spaces. The generalization of the proof-theoretic analysis itself was ultimately relegated to the forthcoming paper [19].
Now, this work raises the question of whether the introduction of this new abstract notion only provides a unified proof of the theorem for the two known cases or whether one has properly extended the class of spaces for which the theorem holds. This is the question that this paper aims to answer.
Towards that end, we take the existing construction of a product of metric spaces and show that it extends all the way to uniformly smooth hyperbolic spaces. Uniform convexity may possibly not be preserved under products, but we show that a specific property, which is enough for the proof of Reich’s theorem, is. Finally, we give an example of a product space for which we show that it cannot be either a CAT(0) space or a bounded subset of a normed space.

### v2

@@ 1 Introduction
If $X$ is a metric space possessing enough convexity structure, and $T:X\to X$ is a nonexpansive mapping, one can define, for any $x\in X$, the point $x_{t}$, called a ‘resolvent’, by the implicit definition $x_{t}=tTx_{t}+(1-t)x$ (existence and uniqueness being guaranteed by the Banach contraction principle). One may then ask about the strong convergence of the path $(x_{t})$ (for $t\to 1$), and, for $X$ being a bounded convex subset of a Hilbert space, this was first proven in 1967, independently by Browder [5] and Halpern [8]. A little more than ten years later, the corresponding result for uniformly smooth Banach spaces, requiring significantly new ideas, was proven by Reich [20].
@@
The analysis of the Browder-Halpern result was obtained by Kohlenbach in 2011 [11]; its sequel, the work on Reich’s theorem, had to wait, like the original theorem itself, a decade until it was finally achieved by Kohlenbach and the author [14], being one of the most intricate results ever achieved by the program of proof mining (see also the author’s recent extension [23] to Reich’s original setting of accretive operators). The goals of these endeavours were to obtain quantitative results, namely so-called ‘rates of metastability’; what is more relevant here is that a proof-theoretic analysis usually strips down a proof to its bare essence, leaving it more suitable to generalization.
Such a generalization was needed, as per Pinto [18], since the theorem had also been proven, in the meantime, in the setting of CAT(0) spaces by Saejung [21] (and then proof-theoretically analyzed by Kohlenbach and Leuştean [13]), and, thus, one had to have a framework to encompass both CAT(0) spaces and uniformly smooth Banach spaces, this being achieved by the class of uniformly smooth hyperbolic spaces. One additional issue was that Kohlenbach and the author, in order to tame the proof of Reich’s theorem and make it amenable to analysis and to generalization, introduced the additional uniform convexity hypothesis on the Banach space (thus, still maintaining $L^{p}$ spaces as a particular case of the theorem) – fortunately, this notion had already been formulated in the hyperbolic setting by Leuştean [16, 17] in the form of $UCW$-hyperbolic spaces. Thus, Pinto was able to show that Reich’s theorem and its consequences do hold for uniformly smooth $UCW$-hyperbolic spaces. The generalization of the proof-theoretic analysis itself was ultimately relegated to the forthcoming paper [19].
Now, the introduction of this new abstract notion raises the question of whether it solely provides a unified proof of the theorem for the two known cases or whether one has properly extended the class of spaces for which the theorem holds. This is the question that this paper aims to answer.
Towards that end, we take the existing construction of a product of metric spaces and show that it extends all the way to uniformly smooth hyperbolic spaces. Uniform convexity may possibly not be preserved under products, but we show that a specific property, which is enough for the proof of Reich’s theorem, is. Finally, we give an example of a product space for which we show that it cannot be either a CAT(0) space or a convex subset of a normed space.